If a regular octagon has a side-length 1, find the distance between opposite sides.
Answer by richwmiller(17219)

if one side is 1 unit
2a^2=1^2
a^2=1/2
a=sqrt(1/2)
distance = 2sqrt(.5)+1
which is twice one leg of the right isosceles triangle with hypotenuse of 1 plus one for the one unit side